OpenStudy (anonymous):

find the indefinite integral : integral(sin(4x)^3 cos(4x)^4)dx

OpenStudy (anonymous):

Start with u = 4x and du = 4x dx

OpenStudy (anonymous):

Now apply the reduction formula.

OpenStudy (anonymous):

Did you get it?

OpenStudy (dumbcow):

wouldn't you want to make the substitution: \[\sin^{3}(4x) = \sin(4x)(1-\cos^{2} (4x))\] \[u = \cos(4x)\] \[du = -4\sin (4x) dx\] \[\rightarrow -\frac{1}{4}\int\limits_{}^{}(1-u^{2})u^{4} du\]

OpenStudy (anonymous):

@dumbcow nice one! Its a lot less messier!

OpenStudy (dumbcow):

thanks sgamam, can you finish from there

OpenStudy (anonymous):

yeah I just finished it thank you both so much for the help!